OpenStudy (anonymous):

what is the solution to x-5+3=6

OpenStudy (anonymous):

x - 2 = 6 +2 + 2 addition property of equality x = 8

OpenStudy (anonymous):

x-5=6-3 x-5=3 x=3+5 x=8